OpenStudy (anonymous):

simplify the following expression. √6(6+√3)

OpenStudy (anonymous):

@RoseDryer please help?

OpenStudy (anonymous):

\[\sqrt{6(6+\sqrt{3}}\] \[\sqrt{6(6+1.7321}\] \[\sqrt{6(7.7321)}\] \[\sqrt{46.3923}\] \[=6.8111\]